OpenStudy (anonymous):

Help anyone? :D thank you <3 y varies jointly with p and q. When p = 7 and q = 2, y = 70. Find y when p = 8 and q = 4. y = 120 y = 160 y = 200 y = 240

OpenStudy (anonymous):

My friend taught me how to do these but i forgot :(

OpenStudy (anonymous):

let y=kpq as y varies jointly with p and q now When p = 7 and q = 2, y = 70 hence k=y/pq=70/(7*2)=5 hence y=5pq p = 8 and q = 4. or y=5*8*4=160 or y = 160

OpenStudy (anonymous):

Thank you <3

OpenStudy (anonymous):

welcome dear